OpenStudy (anonymous):

Round the square root of 77 to the nearest hundredth

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@phi

OpenStudy (satan ):

You could plug it into a calculator if you have one?

OpenStudy (satan ):

Also, \(\sqrt{77}\) is between 8(\(\sqrt{64}\)) and 9(\(\sqrt{81}\))

OpenStudy (anonymous):

Is the answer 8.77

OpenStudy (anonymous):

@Satan 

OpenStudy (anonymous):

I got 8.77

OpenStudy (anonymous):

Yes, and rounded correctly

OpenStudy (anonymous):

Thx

OpenStudy (satan ):

Yea, it's right.